OpenStudy (aleah54):

can someone please help quick Choose the correct simplification of the expression (3x)4 can you explain how you got your answer

OpenStudy (keito_):

(3x)*4 3x*4 3*4x 3*4 =12x

OpenStudy (mrnood):

I assume that you mean \[(3x)^{4}\] is that correct?

OpenStudy (mrnood):

You can see (for example) that \[(2x)^{2} = (2x) * (2x) = 2*2*x*x = 2^{2}*x ^{2}\] So you can put the power 4 into each factor inside the bracket...

OpenStudy (aleah54):

the answer is 81x^4

OpenStudy (aleah54):

how did they get that answer

OpenStudy (aleah54):

@rc3212 @RainbowDash1145 @quickstudent @T-Dawg02 @mathmale

OpenStudy (mrnood):

I have given you the answer above. more generally \[(ab)^{n} = a ^{n}b ^{n}\] in you case a = 3 b= x n= 4 so use my general case - put in your numbers and your answer will be there...

OpenStudy (mrnood):

Use ^ to represent powers if oyu do not know the equation editor...

OpenStudy (mrnood):